OpenStudy (anonymous):

The formula for finding the present value of an item that depreciates yearly is v=c-crt. In this formula, v is the present value, c is the original cost, r is the rate of depreciation per year, and t is the number of years that have passed. After 1 years, what is the value of a car originally costing 26,000 that depreciated at a rate of 0.1 per year?

OpenStudy (anonymous):

what is rt?

OpenStudy (anonymous):

I guess the rate

OpenStudy (anonymous):

The rate r is depreciation per year, which is given as __ The time t is 1 year.

OpenStudy (anonymous):

That's exactly how the problem is written.

OpenStudy (anonymous):

I think you can solve this problem, if I help you. I know I can solve it.

OpenStudy (anonymous):

I don't know how to start it

OpenStudy (anonymous):

I mean I tried to plug things in though

OpenStudy (anonymous):

The rate r is depreciation per year, which is given as __ is that a percent, or a decimal? Usually it would be given as a percent

OpenStudy (anonymous):

Let's assume it's 10% per year, r = 0.1 OK?

OpenStudy (anonymous):

How many years does the problem ask you to calculate ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

After 1 years, what is the value of a car originally costing 26,000 that depreciated at a rate of 0.1 per year?

OpenStudy (anonymous):

So how many years is that?

OpenStudy (anonymous):

10%?

OpenStudy (anonymous):

1 year.

OpenStudy (anonymous):

If r = 0.1 and t = 1, what is rt = ?

OpenStudy (anonymous):

So now I'd like to factor the equation you were given v=c-crt v = c(1 - rt) We just found that rt = 0.1. c = 26000 can you do the rest?

OpenStudy (anonymous):

23400 is answer

OpenStudy (anonymous):

Sounds correct to me.

OpenStudy (anonymous):

Thank you !